OpenStudy (anonymous):

Last week, Eva spent 8 3/8 hours on homework. This week, she spent 10 5/8 hours. How many more hours did she spend this week?

OpenStudy (bahrom7893):

10 5/8 - 8 3/8

OpenStudy (anonymous):

10 5/8-8 3/8=2 1/4

OpenStudy (bahrom7893):

good job jm92

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

haha it worked

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{T}}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{n}}\color{blue}{\normalsize\text{k}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{U}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{P}}\color{#9c9a2e}{\normalsize\text{e}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{p}}\color{purple}{\normalsize\text{l}}\color{purple}{\normalsize\text{e}}\color{red}{\normalsize\text{!}}\color{orange}{\normalsize\text{!}}\color{#9c9a2e}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{T}}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{n}}\color{blue}{\normalsize\text{k}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{U}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{P}}\color{#9c9a2e}{\normalsize\text{e}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{p}}\color{purple}{\normalsize\text{l}}\color{purple}{\normalsize\text{e}}\color{red}{\normalsize\text{!}}\color{orange}{\normalsize\text{!}}\color{#9c9a2e}{\normalsize\text{}}\end{array}